==========some issues have been fixed====================== 1) Fixed left arm bcm d10. It was unplugged downstairs previously. 2) Found reason why pileup E4NA (tag & electron narrow grp 4) has no readings. It turned out the channel in NIM-ECL does not work. We copied this signal using a module 758 and send it to the broken NIM-ECL channel (NIM-ECL 5, chan15) and NIM-ECL 6, chan 8.
Found a happex run (12348) after these been done. Can see bcm d10 of left arm and pileup E4NA (NIM-ECL6, chan 8) from raw data (still zero for pileup E4NA in old channel).
3) Also found E4WA no readings in raw data while looking at run 12348. Do not quite remember what we did, but found signal back afterward.
============sync signal added==================
Luis showed us where the sync1 and sync2 signals are in both arms. We fan
them out and send copies to PVDIS DAQ. These signals are in NIM-ECL 6,
chan 9 (gate/synch1, scaler9_8) and chan10 (ungate/synch2, scaler9_9) for
right arm; the last NIM-ECL (starting from left, scaler13_), chan 9
(gate/synch1) and chan10 (ungate/synch2) in left arm.
A copy of this log entry has been emailed to: rom@jlab.org, xiaochao@jlab.org, rsubedi@jlab.org, pkpan@jlab.org, sarahp@jlab.org, dwang@jlab.org, deng@jlab.org